# On Topological Groups of Automorphisms

### v1

@@ 1. Introduction
The focus of this study is the set of homeomorphic bijections (automorphisms) of a space $X$ with itself, denoted by $Hom(X)$, endowed with the topology of point-wise convergence. The resulting space is denoted by $Hom_{p}(X)$. Recall that a standard basic set of this topology is in the form $\{h\in Hom(X):h(x_{i})\in O_{i},i=1,...,n\}$, where $x_{1},...,x_{n}$ are arbitrary fixed elements of $X$ and $O_{1},...,O_{n}$ are arbitrary fixed open subsets of $X$. It is known that $Hom(X)$ is an algebraic group with respect to the operation of composition, which is always assumed to be the group operation in this study. It is well-known that $Hom_{p}(X)$ need not be a topological group. For example, $Hom_{p}(\mathbb{R}^{2})$ is not a topological group as mentioned in one of the exercises in [1]. There are also many examples of spaces for which the respective structure is a topological group. For example, $Hom_{p}(L)$ is a topological group for any connected linearly ordered space $L$ ([4]). Moreover, it was shown in [2] that $Hom_{p}(X)$ is a topological group whenever $X$ is the union of finitely many closed connected linearly ordered subspaces. In this study, will will focus on finding spaces $X$ for which $Hom_{p}(X)$ is a topological group. To justify a clear lack of homogeneity in our candidates, we first start by proving that too much of homogeneity in $X$ is often the reason for $Hom_{p}(X)$ not being a topological group. Clearly, $Hom_{p}(X)$ is a topological group if $X$ has no automorphisms except the identity. The space in this study are rich with automorphisms.
@@ 2. Ones That Are Not
For the remainder of this section, we say that $X\subset Y$ is $h$-embedded in $Y$ if any homeomorphisim $h$ from $X$ onto $X$ can be extended to a homeomorphism from $Y$ onto $Y$.
@@ 3. Ones That Are
Let $X$ be a space and $x,y\in X$. We say that $x$ is equivalent to $y$ if $f(x)=y$ for some $f\in Hom(X)$. We say that $X$ is locally unique at $x$ if there exists an open neighborhood $O$ of $x$ such that $x$ is not equivalent to any $y$ in $O\setminus\{x\}$. Loosely speaking, all look-alikes of $x$ are very far from $x$. There are many such spaces among classical examples. In particular, any subspace of an ordinal is such. Moreover, any subspace of $\alpha^{n}$ is such for any ordinal $\alpha$ and any positive integer $n$. Observe that if $x\in X$ is not equivalent to any element in $O\setminus\{x\}$, then $f(x)$ is not equivalent to any element in $f(O)\setminus\{f(x)\}$ for any $f\in Hom(X)$.
@@
Let $\mathcal{T}_{p}$ and $\mathcal{T}_{d}$ be the topologies of $Hom_{p}(X)$ and $Hom_{d}(X)$, respectively. By our earlier discussion, it suffices to show that $\mathcal{T}_{d}\subset\mathcal{T}_{p}$. For this, fix any open neighborhood $U$ of $f$ in $Hom_{p}(X)$. There exist $x_{1},...,x_{n}\in X$ and open sets $O_{1},...,O_{n}\subset X$ such that $f\in V=\{h\in Hom(X):h(x_{i})\in O_{i},i=1,...,n\}\subset U$. We may asume that each $O_{i}$ is small enough so that $f(x_{i})$ is not equivalent to any element in $O_{i}\setminus\{f(x_{i})\}$. Therefore, $h(x_{i})=f(x_{i})$ for each $i=1,...,n$ and each $h\in V$. Hence $V=\{h\in Hom(X):h(x_{i})=f(x_{i}),i=1,...,n\}$, which completes the proof. ∎
Corollary 3.5.
Let $X$ be a subspace of the product of finitely many ordinals. Then $Hom_{p}(X)$ is a topological group.
Question 3.6.
Let $X$ be a scattered space. Is $X$ locally unique? Is $Hom_{p}(X)$ a topological group?

### v2

@@ 1. Introduction
The focus of this study is the set of homeomorphic bijections (automorphisms) of a space $X$ with itself, denoted by $Hom(X)$, endowed with the topology of point-wise convergence. The resulting space is denoted by $Hom_{p}(X)$. Recall that a standard basic set of this topology is in the form $\{h\in Hom(X):h(x_{i})\in O_{i},i=1,...,n\}$, where $x_{1},...,x_{n}$ are arbitrary fixed elements of $X$ and $O_{1},...,O_{n}$ are arbitrary fixed open subsets of $X$. It is known that $Hom(X)$ is an algebraic group with respect to the operation of composition, which is always assumed to be the group operation in this study. It is well-known that $Hom_{p}(X)$ need not be a topological group. For example, $Hom_{p}(\mathbb{R}^{2})$ is not a topological group as mentioned in one of the exercises in [1]. There are also many examples of spaces for which the respective structure is a topological group. For example, $Hom_{p}(L)$ is a topological group for any connected linearly ordered space $L$ ([5]). Moreover, it was shown in [2] that $Hom_{p}(X)$ is a topological group whenever $X$ is the union of finitely many closed connected linearly ordered subspaces. In this study, will will focus on finding spaces $X$ for which $Hom_{p}(X)$ is a topological group. To justify a clear lack of homogeneity in our candidates, we first start by proving that too much of homogeneity in $X$ is often the reason for $Hom_{p}(X)$ not being a topological group. Clearly, $Hom_{p}(X)$ is a topological group if $X$ has no automorphisms except the identity. The spaces in this study are rich with automorphisms.
@@ 2. Ones That Are Not
In Corollary 2.3 and the preceding lemmas, we cannot replace ”strong $n$-homogeneity” with ”$n$-homogeneity”. Recall that $X$ is $n$-homogeneous if given any two $n$-sized subsets $A$ and $B$ of $X$ there exists $h\in Hom(X)$ such that $f(A)=B$ (see [6] for references). Note that $\mathbb{R}$ is $n$-homogeneous for any positive $n$ but not strongly $n$-homogeneous for $n>2$. Therefore, $\mathbb{R}$ is a witness that ”strongly” cannot simply be dropped from the mentioned statements.
For the remainder of this study, we say that $X\subset Y$ is $h$-embedded in $Y$ if any homeomorphisim $h$ from $X$ onto $X$ can be extended to a homeomorphism from $Y$ onto $Y$.
@@ 3. Ones That Are
Let $X$ be a space and $x,y\in X$. We say that $x$ is equivalent to $y$ if $f(x)=y$ for some $f\in Hom(X)$. We say that $X$ is locally unique at $x$ if there exists an open neighborhood $O$ of $x$ such that $x$ is not equivalent to any $y$ in $O\setminus\{x\}$. Loosely speaking, all look-alikes of $x$ are very far from $x$. There are many such spaces among classical examples. In particular, any subspace of an ordinal is such. Moreover, any subspace of $\alpha^{n}$ is such for any ordinal $\alpha$ and any positive integer $n$. It is also observed in [4] that scattered spaces are such as well. Observe that if $x\in X$ is not equivalent to any element in $O\setminus\{x\}$, then $f(x)$ is not equivalent to any element in $f(O)\setminus\{f(x)\}$ for any $f\in Hom(X)$.
@@
Let $\mathcal{T}_{p}$ and $\mathcal{T}_{d}$ be the topologies of $Hom_{p}(X)$ and $Hom_{d}(X)$, respectively. By our earlier discussion, it suffices to show that $\mathcal{T}_{d}\subset\mathcal{T}_{p}$. For this, fix any open neighborhood $U$ of $f$ in $Hom_{p}(X)$. There exist $x_{1},...,x_{n}\in X$ and open sets $O_{1},...,O_{n}\subset X$ such that $f\in V=\{h\in Hom(X):h(x_{i})\in O_{i},i=1,...,n\}\subset U$. We may assume that each $O_{i}$ is small enough so that $f(x_{i})$ is not equivalent to any element in $O_{i}\setminus\{f(x_{i})\}$. Therefore, $h(x_{i})=f(x_{i})$ for each $i=1,...,n$ and each $h\in V$. Hence $V=\{h\in Hom(X):h(x_{i})=f(x_{i}),i=1,...,n\}$, which completes the proof. ∎
In [4], the author proves that $Hom_{p}(X)=Hom_{d}(X)$ for scattered spaces and concludes that $Hom_{p}(X)$ is a topological group. Even though Lemma 3.4 does not add any new interesting spaces to our collection, we included it to double down on our original claim that we have to sacrifice strong homogeneity properties in search for spaces $X$ for which $Hom_{p}(X)$ is a topological group.